# A Study on Adversarial Robustness of Discriminative Prototypical Learning

### v1

@@ 4.4 Robustness on Common Corruptions
Method Avg Acc (%)
AT 76.16
AT_HE 73.20
TRADES 73.84
MART 69.04
Adv-DPNP 76.73
[…]
Figure 3: Comparison of models’ performance against different severity levels of the CIFAR-10-C dataset. Each bar represents an average of 19 corruption types for a given level of corruption strength.
We evaluate accuracy of the adversarially trained models from Section 4.1 on the CIFAR-10-C dataset without any additional fine-tuning. Figure 3 illustrates the accuracy for each level of severity averaged across all corruptions. Adv-DPNP consistently improves accuracy at each severity level and outperforms other methods. In addition, following [41], we report the average accuracy across all severities in Table 6. Adv-DPNP increases the average accuracy by 0.57%, with AT in second place. These results demonstrate that Adv-DPNP can effectively improve model performance under common real-world corruptions, making it an effective approach for improving robustness.
@@ 4.5 Verifying the Absence of Gradient Obfuscation
In this section, we follow the guidelines outlined in [42, 43] to verify the absence of gradient obfuscation in our Adv-DPNP framework. For this purpose, we evaluate the Adv-DPNP model trained as described in Section 4.1 by performing a series of tests under different conditions.
@@
[…]
Increasing Number of Iterations and Random Starts: To further investigate the potential for gradient obfuscation, we report the accuracy and CE loss values when increasing the number of iterations and applying random starts for PGD attacks for CIFAR-10 and CIFAR-100, as shown in Tables 8 and 8. The results indicate that both accuracy and loss values change marginally with increases in the number of iterations and random starts. This behavior confirms that the iterative PGD attack converges correctly.
Adaptive Attacks. We also evaluate the robustness of the Adv-DPNP model by generating adversarial examples using our composite Adv-DPNP loss (Equation 7) objectives for PGD attacks.
@@
[…]
Table 9 compares the model accuracy on CIFAR-10 and CIFAR-100 under CE and our composite loss. The results show only a negligible difference in accuracy between the two loss objectives, which confirms that the robustness of Adv-DPNP is not due to its loss surface making optimization difficult for attacks with CE [42].
In conclusion, the results of these experiments indicate that Adv-DPNP does not exhibit gradient obfuscation. Therefore, the Adv-DPNP model is robust against various adversarial attacks, and our evaluations are reliable.
@@ 5 Conclusions
In this work, we proposed Adv-DPNP, an adversarial training framework that effectively integrates discriminative prototype-based learning to enhance both adversarial robustness and clean data accuracy. By unifying class prototypes with classifier weights, and employing a dual-branch training mechanism that decouples prototype updates from adversarial data, Adv-DPNP not only improves model performance under adversarial attacks but also better preserves the original prototypes. In addition, benefiting from discriminative terms in its composite loss function, our proposed method is able to offer an organized latent space, leading to more robust representations.
We demonstrated that Adv-DPNP outperforms state-of-the-art adversarial defenses on benchmark datasets such as CIFAR-10 and CIFAR-100, achieving significant improvements in both clean accuracy and robustness against a wide range of adversarial perturbations. Additionally, we showed that Adv-DPNP provides superior performance under common corruptions on the CIFAR-10-C dataset. Through extensive experiments and analysis of feature space discrimination, we validated that Adv-DPNP maintains a well-structured feature space with high intra-class compactness and inter-class separation, even when adversarial perturbations are applied. Furthermore, our analysis confirmed that Adv-DPNP does not suffer from gradient obfuscation, ensuring the stability and effectiveness of the model against various attack strategies.

### v2

@@ 4.4 Robustness against Common Corruptions
To assess the natural robustness of our method, we evaluate the performance of adversarially trained models on the original CIFAR-10 dataset using common corruptions dataset CIFAR-10-C for test [16]. CIFAR-10-C is created by corrupting the original CIFAR-10 test images with 19 corruptions. These corruptions are categorized into four types including noise, blur, weather, and digital, with each type having five levels of severity or intensity, which increases from level one to five.
We evaluate the accuracy of the adversarially trained models from Section 4.1 on the CIFAR-10-C dataset without any additional fine-tuning. Figure 3 illustrates the accuracy for each level of severity averaged across all corruptions. Adv-DPNP consistently improves accuracy at each severity level and outperforms other methods. In addition, following [16], we report the average accuracy across all severities in Table 5. Adv-DPNP increases the average accuracy by 0.57%, with AT in second place. These results demonstrate that Adv-DPNP can effectively improve model performance under common real-world corruptions, making it an effective approach for improving robustness.
@@ 4.5 Verifying the Absence of Gradient Obfuscation
In this section, we follow the guidelines outlined in [4, 1] to verify the absence of gradient obfuscation in our Adv-DPNP model. For this purpose, we evaluate Adv-DPNP with the ResNet-18, trained according to Section 4.1, by performing a series of tests under different conditions.
@@
Increasing Number of Iterations and Random Starts: To further investigate the potential for gradient obfuscation, we report the accuracy and CE loss values when increasing the number of iterations and applying random starts for PGD attacks for CIFAR-10 and CIFAR-100, as shown in Tables 7 and 7. The results indicate that both accuracy and loss values change marginally with increases in the number of iterations and random starts. This behavior confirms that the iterative PGD attack converges correctly.
@@
Adaptive Attacks. We also evaluate the robustness of the Adv-DPNP model by generating adversarial examples using our composite Adv-DPNP loss (Equation 7) objectives for PGD attacks. Table 8 compares the model accuracy on CIFAR-10 and CIFAR-100 under CE and our composite loss. The results show only a negligible difference in accuracy between the two loss objectives, which confirms that the robustness of Adv-DPNP is not due to its loss surface making optimization difficult for attacks with CE [4]. In conclusion, the results of these experiments indicate that Adv-DPNP does not exhibit gradient obfuscation. Therefore, the Adv-DPNP model is truly robust and our evaluations are reliable.
@@ 5 Conclusions
In this work, we proposed Adv-DPNP, an adversarial training framework that effectively incorporates discriminative prototype-based learning to enhance both adversarial robustness and clean data accuracy. By unifying class prototypes with classifier weights, and employing a dual-branch training procedure that decouples prototype updates from adversarial data, Adv-DPNP not only obtains robust representations which improve model performance under adversarial attacks but also better preserves clean accuracy and the original prototypes. In addition, benefiting from discriminative terms in its composite loss function, our proposed method is able to offer an organized latent space, leading to better generalization.
We demonstrated that Adv-DPNP outperforms common baseline methods and state-of-the-art adversarial defenses on benchmark datasets such as CIFAR-10, CIFAR-100, and SVHN, achieving significant improvements in both clean accuracy and robustness against a wide range of adversarial perturbations. Additionally, we showed that Adv-DPNP provides superior performance under common corruptions on the CIFAR-10-C dataset. Through extensive experiments and analysis of feature space discrimination, we validated that Adv-DPNP maintains a well-structured feature space with high intra-class compactness and inter-class separation, even when strong adversarial perturbations are applied. Furthermore, our analysis confirmed that our proposed method does not suffer from gradient obfuscation, ensuring the stability and effectiveness of the model against various attack strategies.